英語 (English)

発信型英語 (Communicative English) (工 ((化 A ー 1)2 年))

早内-プリングルジュディス · 非常勤講師 2 単位後期火 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】To give students opportunities to develop their English abilities through the exploration and discussion of a variety of themes. Students will develop their real-life communicative skills while expanding their vocabulary, grammatical ability, reading ability and fluency. Various media forms will be used in this class.

【授業の概要】 Speaking, reading, presentation Notes: 1. Students are expected to speak and write only in English during class. They are expected to work not only in class, but outside of class as well. Homework is required, and they are expected to complete it even if they are absent from class. 2. The following class schedule is tentative and only shows the general contents which will be adjusted based on student input. There may be other modifications based on class size and other factors.

【キーワード】 Communicative English， fluency 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】In today’s globalized society, English has become the common

language of communication and is necessary to gain more knowledge of world events, and process information. This class requires students to put their skills into practice, and provides tools to increase fluency and confidence in using English.

【授業の計画】 1. General Class Guidance Introduction; Survey & Interview

2. Presentations: Basic Skills

3. Topic 1: Pop Culture: Trending now: Explore the topic

4. Topic 1: Discuss

5. Topic 1: Expand

6. Topic 2: Health in the News: Explore the topic

7. Topic 2: Discuss & interview

8. Mid-term Evaluation/Group Presentation

9. Topic 3: Regional Events: What’s happening in our world? Explore the topic 10. Topic 3: Discuss

11. Topic 3: Expand

12. Topic 4: Technology: Innovations that have changed our world: Explore the topic

13. Topic 4: Discuss

14. Topic 4: Expand

15. Final Examination; Individual Presentations 1

16. Individual Presentations 2

【教科書】 Class handouts and other materials will be provided as necessary. Some

content will also be student-generated. 【参考書等】 [参考資料] 【成績評価の方法】 Grading will be as follows: Class Attendance and participation (including homework, etc):40% Mid-term and final examinations: 20% Presentations: Individual & Group: 20% Media & Current Event Journal: 20% 【再試験の有無】 Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221218 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 早内-プリングル (juditheph@yahoo.com) ���� (オフィスアワー: by appoint ment only)

English

Communicative English (工 ((化 A ー 1)2 年))

Judith Hayauchi · Part-time Lecturer 2 units 後期火 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ To give students opportunities to develop their English abilities through the exploration and discussion of a variety of themes. Students will develop their real-life communicative skills while expanding their vocabulary, grammatical ability, reading ability and fluency. Various media forms will be used in this class.

Outline⟩ Speaking, reading, presentation Notes: 1. Students are expected to speak and write only in English during class. They are expected to work not only in class, but outside of class as well. Homework is required, and they are expected to complete it even if they are absent from class. 2. The following class schedule is tentative and only shows the general contents which will be adjusted based on student input. There may be other modifications based on class size and other factors.

�������⟩ Communicative English, fluency Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ In today’s globalized society, English has become the common language of

communication and is necessary to gain more knowledge of world events, and process information. This class requires students to put their skills into practice, and provides tools to increase fluency and confidence in using English.

Schedule⟩

1. General Class Guidance Introduction; Survey & Interview

2. Presentations: Basic Skills

3. Topic 1: Pop Culture: Trending now: Explore the topic

4. Topic 1: Discuss

5. Topic 1: Expand

6. Topic 2: Health in the News: Explore the topic

7. Topic 2: Discuss & interview

8. Mid-term Evaluation/Group Presentation

9. Topic 3: Regional Events: What’s happening in our world? Explore the topic

10. Topic 3: Discuss 11. Topic 3: Expand

12. Topic 4: Technology: Innovations that have changed our world: Explore the topic

13. Topic 4: Discuss

14. Topic 4: Expand

15. Final Examination; Individual Presentations 1

16. Individual Presentations 2

Textbook⟩ Class handouts and other materials will be provided as necessary. Some

content will also be student-generated. Reference⟩ [参考資料] Evaluation Criteria⟩ Grading will be as follows: Class Attendance and participation (including homework, etc):40% Mid-term and final examinations: 20% Presentations: Individual & Group: 20% Media & Current Event Journal: 20%

Re-evaluation⟩ Allowable if the student has attended more than 2/3 of the classes

and has a score of 50% or higher. Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221218 Contact (Office-Hour, Room, E-mail)⟩

⇒ Hayauchi (juditheph@yahoo.com) ���� (Office Hour: by appointment only)